OpenStudy (anonymous):

Solve the equation and identify the shape.. plz help if you can 7y^2+x^2=64 x+y=4

OpenStudy (dumbcow):

looks like an ellipse and a line

myininaya (myininaya):

the shape of each equation? the first one is an ellipse since both the coefficient of x^2 and y^2 have sign same but have differenet coefficient and yeah second one is a line go cow! :)

OpenStudy (dumbcow):

what are we solving for? or do you need them in standard form

myininaya (myininaya):

\[x=4-y => 7y^2+(4-y)^2=64\]

myininaya (myininaya):

\[7y^2+16-8y+y^2=64\]

myininaya (myininaya):

\[8y^2-8y+16-64=0\]

myininaya (myininaya):

\[8y^2-8y-48=0\] \[y^2-y-6=0\]

OpenStudy (anonymous):

its supposed to be like this